40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 19 10:21:54 2017, Ashley, Update, General, Preliminary Microphone Data xend_psd.pngxblrms.pngIMG_3734.JPG
    Reply  Mon Feb 13 17:39:45 2017, Ashley, Update, General, Preliminary Microphone Data Update figure_1.pngx_and_y_blrms_03.pngxblrms_median.png
Message ID: 12826     Entry time: Mon Feb 13 17:39:45 2017     In reply to: 12738
Author: Ashley 
Type: Update 
Category: General 
Subject: Preliminary Microphone Data Update 
  • Problems that have occurred since my last post: All of the sudden, I was getting very strange data that was very quiet and did not match the previous input range of my last locations (see attachment). After resoldering the custom bnc connection cables with Lydia, which were in disrepair, and checking almost everything we could think of, we found that the gain dial on the preamp was turned all the down. Immediately after it was fixed, the data returned to expected values (based on neighboring locations and data taken at the last location before the problem occurred). 
  • Updates: Since my last post, I have created a normalized blrms color map in addition to the one I already have. Additionally, I have started working on plotting the color maps next to a labeled, to-scale drawing of the lab, but have yet to complete it. 
  • Attachment 1: comparison of the psds
  • Attachment 2: blrms color map
  • Attachment 3: normalized color map

Brief Summary: I am currently looking at the acoustic noise around both arms to see if there are any frequencies from machinery around the lab that stand out and to see what we can remove/change. I am using a Bluebird microphone suspended with surgical tubing from the cable trays to isolate it from vibrations. I am also using a preamp and the SR875 spectrum analyzer taking 6 sets of data every 1.5 meters (0 to 200Hz, 200Hz to 400Hz, 400z to 800Hz, 800Hz to 3200Hz, 3.2kHz to 12kHz, 12kHz to 100kHz).


·                Attachment 1 is a PSD of the first 3 measurements (from 0 to 12kHz) that I took every 1.5 meters along the x arm with the preamp and spectrum analyzer

·                Attachment 2 is a blrms color map of the first 6 sets of data I took (from 2.4m to 9.9m) 

·                Attachmetn 3 is a picture of the microphone set up with the surgical tubing 

Problems that occurred: settings on the preamp made the first set of data I took significantly smaller than the data I took with the 0dB button off and the last problem I had was the spectrum analyzer reading only from -50 to -50 dBVpk




Attachment 1: figure_1.png  46 kB  | Hide | Hide all | Show all
Attachment 2: x_and_y_blrms_03.png  24 kB  | Show | Hide all | Show all
Attachment 3: xblrms_median.png  29 kB  | Show | Hide all | Show all
ELOG V3.1.3-